# SCAIL: Towards Studio-Grade Character Animation via In-Context Learning of 3D-Consistent Pose Representations

### v1

@@ 4.1 Implementation Details
We train two versions (1.3B and 14B) of our model. The 1.3B model is finetuned from the Wan2.1-1.3B-Fun-Inp backbone on our pretraining dataset for 6,000 steps with a batch size of 96 and a learning rate of 1e-5, using 32 NVIDIA H100 GPUs for approximately two days. For the larger 14B model, we finetune it from the Wan2.1-I2V-14B backbone in two stages: during the pretraining stage, we train for 8,000 steps with a batch size of 96 at a learning rate of 1e-5; after convergence, we perform an additional finetuning stage for 400 steps with the same batch size and a reduced learning rate of 4e-6. Training of the 14B model is conducted on 128 NVIDIA H100 GPUs for over four days with sequence parallelism enabled. All models are optimized using AdamW [adamw]. During inference, we set the classifier-free guidance (CFG) scale [cfg] to 4, offering a favorable balance between pose following and video fidelity.
@@ 4.2 Quantitative Evaluation
We conduct a quantitative comparison with state-of-the-art methods on our proposed Studio-Bench. For self-driven part we employed several widely-used quantitative metrics, including PSNR [psnr], SSIM [ssim], LPIPS [lpips], and FVD [fvd]. To evaluate the generated results in the second subset, we design four metrics: (1) Motion Accuracy, which measures how faithfully the generated motion follows the driving signal in a frame-by-frame manner. (2) Kinesiology Consistency, evaluating whether joint rotations and body movements remain anatomically feasible and temporally coherent, penalizing sudden twists or physically impossible poses that break natural motion continuity. (3) Physical Consistency, assessing whether the generated motions comply with basic physical constraints such as gravity, support, and momentum conservation, penalizing unrealistic behaviors like hovering in midair. (4) Identity Similarity, measuring the consistency of the subject’s appearance with the reference image. We convey a detailed blinded user study to collect these metrics, letting users vote for the best in baseline comparison. We also conduct a Win/Tie/Lose study to evaluate our performance between two commonly used open-source frameworks, as well as the closed-source commercial product Viggle [viggle], omitting UniAnimate-DiT[Unianimate-dit] due to lack of competitiveness under studio-grade demandings. Viggle is widely believed to rely on a 3D foundation model rather than video diffusion and can be a strong baseline for frame-by-frame motion accuracy under complex scenarios.
@@ 4.4 Ablation Study
We conduct ablation studies on SCAIL-1.3B model to evaluate the contribution of each component. For reference, we also compare against the U-Net based MimicMotion [MimicMotion], highlighting the advantage of our DiT-based architecture. In the main paper we mainly focus on ablations on model architecture, and more ablations on the pose condition would be provided in Appendix A2.
@@ 5 Conclusion
Rendering Details. To preserve spatial relationships between joints, we represent the human skeleton using cylindrical structures. For efficient rendering of multiple skeletons, we employ a 3D rendering pipeline that first converts the cylinders into spatial voxels, followed by ray marching implemented via [taichi]. This strategy is highly optimized for modern GPUs, introducing negligible computational overhead. To facilitate person discrimination in multi-person motion sequences, we assign distinct color schemes to each individual. This enables the model to directly learn how to distinguish characters from the representation. We observe that this design effectively alleviates identity switching when the positions of the people interchange.
Augmentation Details. As our augmentation strategy is designed to maximally preserve the original motion, we use a high augmentation rate of 0.8 to achieve the balance between the pose-following accuracy and motion transfer robustness. During training, the overlaid 2D hand and face keypoints extracted by DWPose [DWPose] will go through additional augmenation after the 3D adaptation process where 2D keypoints are shifted to match the reformed 3D skeleton after body rescaling and camera manipulation. This step is to minimize the unintended influence of the 2D facial and hand signals on the 3D pose representation.
Performance Tradeoff. For the injection of the representation in full-context settings, we observe that the rendered pose sequence is relatively sparse as most frame areas consist of non-informative black pixels. In our method, spatial pooling to pose sequence can serve as a simple workaround to effectively reduce the contextual pose tokens to $1/4$ and preserves accurate pose-following capability. In addition, full-context pose injection introduces no new parameters except the additional patchify layer to the original model, offering a more streamlined architecture compared to stacking additional DiT layers in residual context-tuning methods [VACE]. In general, considering the significant motion error reduction, we conclude that the modest efficiency trade-off is acceptable, particularly in studio-grade scenarios which prioritize stringent accuracy and stability.
@@
To validate the effectiveness of our proposed 3D-consistent pose representation, we compare with a 2D keypoints-based pipeline (extracted using DWPose [DWPose] and denoted as w/ 2D Pose). Specifically, we implement the 2D augmentation strategy as close as to the 3D version, with similar figure settings and same augmentation ratio of 0.8. Faces and hands are also augmented identically to the 3D version and retargeting logic from [Unianimate-dit] are applied during inference. All model architectures keep the same using full-context pose injection and P-RoPE. As shown in Table A1, the estimation noise and the 2D-3D mismatch noise introduced by the pipeline significantly undermine the model’s performance in the self-driven subset of our challenging Studio-Bench, which involves a high ratio of complex motions.
For the cross-driven subset, 2D pose can easily lead to distorted limbs in generation, as seen in case (1) of Figure A1. Case (2) indicates that when model needs to distinguish the front and back of limbs, the inherent ambiguity of 2D pose can result in incorrect pose interpretations. Qualitative results from the cross-driven subset and quantitative results from the self-driven subset together demonstrate the overall effectiveness of our proposed 3D-based solution. Figure A2 can also demonstrate the unreasonable scaling factors introduced by 2D-based retargeting process during the inference step, which we will discuss in the next section of retarget ablations.

### v2

@@ 4.1 Implementation Details
We train two versions (1.3B and 14B) of our model. The 1.3B model is finetuned from the Wan2.1-1.3B-Fun-Inp backbone on our pretraining dataset for 6,000 steps with a batch size of 96 and a learning rate of 1e-5, using 32 NVIDIA H100 GPUs for approximately two days. For the larger 14B model, we finetune it from the Wan2.1-I2V-14B backbone in two stages: during the pretraining stage, we train for 8,000 steps with a batch size of 96 at a learning rate of 1e-5; after convergence, we perform an additional finetuning stage for 400 steps with the same batch size and a reduced learning rate of 4e-6. Training of the 14B model is conducted on 128 NVIDIA H100 GPUs for over four days with sequence parallelism enabled. All models are optimized using AdamW [21]. During inference, we set the classifier-free guidance (CFG) scale [13] to 4, offering a favorable balance between pose following and video fidelity.
@@ 4.2 Quantitative Evaluation
We conduct a quantitative comparison with state-of-the-art methods on our proposed Studio-Bench. For self-driven part we employed several widely-used quantitative metrics, including PSNR [14], SSIM [40], LPIPS [48], and FVD [32]. To evaluate the generated results in the second subset, we design four metrics: (1) Motion Accuracy, which measures how faithfully the generated motion follows the driving signal in a frame-by-frame manner. (2) Kinesiology Consistency, evaluating whether joint rotations and body movements remain anatomically feasible and temporally coherent, penalizing sudden twists or physically impossible poses that break natural motion continuity. (3) Physical Consistency, assessing whether the generated motions comply with basic physical constraints such as gravity, support, and momentum conservation, penalizing unrealistic behaviors like hovering in midair. (4) Identity Similarity, measuring the consistency of the subject’s appearance with the reference image. We convey a detailed blinded user study to collect these metrics, letting users vote for the best in baseline comparison. We also conduct a Win/Tie/Lose study to evaluate our performance between two commonly used open-source frameworks, as well as the closed-source commercial product Viggle [35], omitting UniAnimate-DiT[38] due to obvious artifacts under studio-grade demandings. Viggle is widely believed to rely on a 3D foundation model rather than video diffusion and can be a strong baseline for frame-by-frame motion accuracy under complex scenarios.
@@ 4.4 Ablation Study
We conduct ablation studies on SCAIL-1.3B model to evaluate the contribution of each component. For reference, we also compare against the U-Net based MimicMotion [49], highlighting the advantage of our DiT-based architecture. In the main paper we mainly focus on ablations on model architecture, and more ablations on the pose condition would be provided in Appendix A2.
@@ 5 Conclusion
Rendering Details. To preserve spatial relationships between joints, we represent the human skeleton using cylindrical structures. For efficient rendering of multiple skeletons, we employ a 3D rendering pipeline that first converts the cylinders into spatial voxels, followed by ray marching implemented via [16]. This strategy is highly optimized for modern GPUs, introducing negligible computational overhead. To facilitate person discrimination in multi-person motion sequences, we assign distinct color schemes to each individual. This enables the model to directly learn how to distinguish characters from the representation. We observe that this design effectively alleviates identity switching when the positions of the people interchange.
Augmentation Details. As our augmentation strategy is designed to maximally preserve the original motion, we use a high augmentation rate of 0.8 to achieve the balance between the pose-following accuracy and motion transfer robustness. During training, the overlaid 2D hand and face keypoints extracted by DWPose [46] will go through additional augmenation after the 3D adaptation process where 2D keypoints are shifted to match the reformed 3D skeleton after body rescaling and camera manipulation. This step is to minimize the unintended influence of the 2D facial and hand signals on the 3D pose representation.
Performance Tradeoff. For the injection of the representation in full-context settings, we observe that the rendered pose sequence is relatively sparse as most frame areas consist of non-informative black pixels. In our method, spatial pooling to pose sequence can serve as a simple workaround to effectively reduce the contextual pose tokens to $1/4$ and preserves accurate pose-following capability. In addition, full-context pose injection introduces no new parameters except the additional patchify layer to the original model, offering a more streamlined architecture compared to stacking additional DiT layers in residual context-tuning methods [17]. In general, considering the significant motion error reduction, we conclude that the modest efficiency trade-off is acceptable, particularly in studio-grade scenarios which prioritize stringent accuracy and stability.
@@
To validate the effectiveness of our proposed 3D-consistent pose representation, we compare with a 2D keypoints-based pipeline (extracted using DWPose [46] and denoted as w/ 2D Pose). Specifically, we implement the 2D augmentation strategy as close as to the 3D version, with similar figure settings and same augmentation ratio of 0.8. Faces and hands are also augmented identically to the 3D version and retargeting logic from [38] are applied during inference. All model architectures keep the same using full-context pose injection and P-RoPE. As shown in Table A1, the estimation noise and the 2D-3D mismatch noise introduced by the pipeline significantly undermine the model’s performance in the self-driven subset of our challenging Studio-Bench, which involves a high ratio of complex motions.
For the cross-driven subset, 2D pose can easily lead to distorted limbs in generation especially when the 1.3B model have difficulty transferring motion to a significant different reference image, as seen in case (1) of Figure A1. Case (2) indicates that when model needs to distinguish the front and back of limbs, the inherent ambiguity of 2D pose can result in incorrect pose interpretations. Qualitative results from the cross-driven subset and quantitative results from the self-driven subset together demonstrate the overall effectiveness of our proposed 3D-based solution. Figure A2 can also demonstrate the unreasonable scaling factors introduced by 2D-based retargeting process during the inference step, which we will discuss in the next section of retarget ablations.